OpenStudy (anonymous):

sqrt{45} times sqrt{60} sqrt{45} times sqrt{60} @Mathematics

OpenStudy (anonymous):

\[\sqrt{45} \times \sqrt{60}\]

jhonyy9 (jhonyy9):

45x60=sqrt2700 =sqrt(9x3x100)=3x10sqrt3 =30sqrt3